A boat can travel 3 times the distance that a car can travel 5 times the distance with ten gallons of gasoline. if the distance travel is know to be 1680 miles, how far can the boat travel

Answer by ankor@dixie-net.com(22740)

A boat can travel 3 times the distance while a car can travel 5 times the distance with ten gallons of gasoline.
if the distance travel is know to be 1680 miles, how far can the boat travel
:
let x = the multiplier
then
3x = the distance traveled by boat
and
5x = the distance traveled by car
Assume the total distance of both is 1680 mi
3x + 5x = 1680
8x = 1680
x = 1680/8
x = 210 is the multiplier
then
3(210) = 630 mi the distance traveled by the boat on 10 gal of gas
:
:
Check by finding the distance the car can travel
5(210) = 1050 mi traveled by the car
total: 1050 + 630 = 1680